Taylor wants to reduce a card that measured 52 centimeters by 16 centimeters to a fraction of its original size. she wants the perimeter of the reduced card to be 34 centimeters.
Answer by jorel555(1290)

The perimeter of the card now is 2x(52+16)=136
136/34=4
So the dimensions of the card have to be reduced to 1/4 of its' original:
52/4=13
16/4=4
Giving you a card that is 13cm x 4cm ☺☺☺☺